OpenStudy (anonymous):

When one U.S. Dollar is worth one point five two Australian Dollars, which of the following could be used as a conversion factor? nineteen one hundred twenty-fifths one and twelve twenty-fifths one and thirteen twenty-fifths five and one-fiftieth

OpenStudy (jack1):

1 = 1.52 so 0.52 = 52/100 =13/25 so 1 = 1 + 13/25 u good from here dude?

OpenStudy (anonymous):

yeah

OpenStudy (jack1):

sweet, slaters